OpenStudy (anonymous):

Evaluate sec 240° without using a calculator by using ratios in a reference triangle.

OpenStudy (anonymous):

unit circle??

OpenStudy (anonymous):

with outt it

hartnn (hartnn):

ok, u know what is cos 60 ?

OpenStudy (anonymous):

pi/3

hartnn (hartnn):

cos pi/3 = ?

OpenStudy (anonymous):

60

hartnn (hartnn):

lol! 60=pi/3 cos 60 = 0.5 ok ?

OpenStudy (anonymous):

ohh ok

hartnn (hartnn):

sec 240 = 1/ cos 240 cos 240 = cos (180+60) cos (180+x) = -cos x now try

OpenStudy (anonymous):

what should i try im a lilttle confused

hartnn (hartnn):

cos (180+x) = -cos x so cos 240 = -cos 60 = -0.5 got this ?

OpenStudy (anonymous):

i think so

hartnn (hartnn):

now sec 240 =1/ cos 240 = 1/(-0.5) = -2 ok ?

OpenStudy (anonymous):

why did u use cos 60 though?

OpenStudy (anonymous):

60 degrees is the reference angle for 240 degrees

hartnn (hartnn):

because cos 60 is known 60 is standard angle in unit circle

OpenStudy (anonymous):

ok

hartnn (hartnn):

any more doubts ?